My age divided by 4 is 17. What is my age?
Answer by addingup(3677)

x/4 = 17 Multiply both sides times 4
(x/4)*4 = 17*4 On the left 4 dividing and 4 multiplying cancel each other:
x = 17*4 Use your calculator, multiply to get your answer (P.S.: you are older than me!)